Exercise 11-03

Novak Company purchased a new plant asset on April 1, 2020, at a cost of $786,000. It was estimated to have a service life of 20 years and a salvage value of $67,800. Novak’ accounting period is the calendar year

Compute the depreciation for this asset for 2020 and 2021 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

$enter a dollar amount

Entry field with incorrect answer

Depreciation for 2021

$enter a dollar amount

Entry field with incorrect answer

Compute the depreciation for this asset for 2020 and 2021 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

$enter a dollar amount

Entry field with incorrect answer now contains modified data

Depreciation for 2021

$enter a dollar amount

Entry field with incorrect answer
Answer #1
Compute the depreciation for this asset for 2020 and 2021 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

(786000/67800/20)*9/12 = 26933

Depreciation for 2021

(786000-67800/20) = 35910
Compute the depreciation for this asset for 2020 and 2021 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

786000*10%*9/12 = 58950

Depreciation for 2021

19650+786000*90%*10%*9/12 = 72705